t/itemView?assignmentProblemID-98937507 Part A 3.55 g of Ca and 3.65 g of F Express your answers using one decimal place separated by a comma. AI Mass % : Ca, F Submit
media%2F01e%2F01e663d5-7b45-4fd2-aa75-21
Answer #1

A) mass % = (mass of individual element / total mass) x 100

mass % Ca = (3.55 / 7.2) x 100

mass % Ca = 49.3 %

mas % F = 100 - 49.3 = 50.7

mass % Ca , F = 49.3 , 50.7

B) (C3H5)2S = 114.21 g /mol

114.21 g (C3H5)2S contain 32 g S

15.4 g (C3H5)2S contain 15.4 x 32 / 114.21 = 4.31 g

mass of S = 4.31 g